What is system physical decomposition ?

if we add spring and damper in series and parellel
Answer #1

Ans a.) System Decomposition : A method of business analysis that dissects a complex business process to show its individual elements. Functional decomposition is used to facilitate the understanding and management of large and/or complex processes and can be used to help solve problems. Functional decomposition is also used in computer engineering to help with software design.

Functional decomposition in systems engineering refers to the process of defining a system in functional terms, then defining lower-level functions and sequencing relationships from these higher level systems functions.[1] The basic idea is to try to divide a system in such a way that each block of a block diagram can be described without an "and" or "or" in the description.

This exercise forces each part of the system to have a pure function. When a system is designed as pure functions, they can be reused, or replaced. A usual side effect is that the interfaces between blocks become simple and generic. Since the interfaces usually become simple, it is easier to replace a pure function with a related, similar function.

For example, say that one needs to make a stereo system. One might functionally decompose this into speakers, amplifier, a tape deck and a front panel. Later, when a different model needs an audio CD, it can probably fit the same interfaces.

Ans b.) springs in parallel

Let two springs with spring constant k1 and k2 are connected in parallel, then resultant spring cont k will be equal to sum of both that is k = k1+k2

springs in Series

Let two springs with spring constant k1 and k2 are connected in series, then resultant spring cont k will be equal to

(1/k)= (1/k1)+(1/k2)

the same results follows for dampers if we add themin series an parallel